# Parallel Key-Value Cache Fusion for Position Invariant RAG

### v1

@@ 7 Conclusion
This paper presents KV-Fusion, a lightweight training scheme aimed at addressing positional bias and improving robustness of decoder-only models in RAG pipeline. KV-Fusion trains language models to be context-order invariant by extracting and pre-filling KV caches with identical positional information, then training decoder-only models using these caches. The results not only highlight the robustness of KV-Fusion in handling a large number of input passages but also its position-invariant property. Our empirical evaluations on three open-domain datasets indicate that KV-Fusion can improve performance and reliability of the RAG system.
@@ 8 Limitations
Second limitation is that our experiments are limited to single-hop question answering, where multi-step reasoning is not required. For example, datasets like HotpotQA Yang et al. (2018) and MuSiQue Trivedi et al. (2022) require multiple passages to derive answers. This work, however, focuses on single-hop question-answering datasets, making it difficult to assess the impact of KV-fusion in multi-hop datasets.
Third limitation is that this work does not fully explore the use of KV-cache for training LLMs. Recently, training LLMs by conditioning key-value caches has gained attention Sun et al. (2024), though our approach remains underexplored in terms of language modeling. However, we present strong empirical results to solve ‘Lost in the middle’ problem. We hope our work can facilitate future studies on utilizing key-value cache for training LLMs.
@@ A.1.1 Algorithm for KV-Fusion
11:  Update parameters of $\mathcal{D}_{\textrm{t}}$ with repect to $Loss$ via gradient descent
@@
To train $\mathcal{D}_{\textrm{t}}$ seaminglessly with huggingface transformersWolf et al. (2020) and PyTorchPaszke et al. (2017), extracted KV-cache need to be reshaped to prefill the caches in $\mathcal{D}_{\textrm{t}}$. To this end, we implement res function down below, which can also process batch of instances.
def reshape_key_value_batches(cur_past_key_values, n_psgs):     """     Reshapes key-value pairs in batches     """     new_key_cache = []     # Iterate through each key-value pair     for k, v in cur_past_key_values:         # Split keys and values into splits (split by instance)         k_splits, v_splits = torch.split(k, n_psgs, dim=0), torch.split(v, n_psgs, dim=0)         # Reshape and concatenate splits (reshape by instance)         k_re = torch.cat([torch.cat(torch.split(k_val, 1, dim=0), dim=2) for k_val in k_splits], dim=0)         v_re = torch.cat([torch.cat(torch.split(v_val, 1, dim=0), dim=2) for v_val in v_splits], dim=0)         # Append processed key-value pair         new_key_cache.append((k_re, v_re))     return tuple(new_key_cache)
@@ A.2.2 Dataset Statistics
NQ and TriviaQA are filtered versions provided by Karpukhin et al. 2020 under the CC BY-NC 4.0. We also use POPQA, available from the HuggingFace datasets [^3] https://huggingface.co/datasets/akariasai/PopQA under MIT License. These datasets are English open domain question answering datasets based on December 2018 Wikipedia snapshot, preprocessed by Karpukhin et al. 2020.
@@ A.5 Hyperparameters for training
Table 5: Across all datasets, we utilize four A100 80 GB GPUs, with a batch size of 2 per device and a gradient accumulation of 8, consuming approximately 12 hours (48 GPU hours). The number of training passages is 20 consisting of one gold context and 19 negative contexts. Contexts are tokenized with a maximum length of 192 tokens using left-padding; if a context exceeds this limit, tokens are truncated from the left. A random sample of 10,000 contexts from the NQ training set showed that 99.0% of contexts fit within this token limit.(Avg: 145 tokens, Std: 14 tokens) The maximum learning rate is set to $2\times 10^{-5}$, using a linear warmup and cosine decay. The warmup ratio is set to 5%. The AdamW optimizer(paged_adamw_32bit) is used with $\beta_{1}=0.9$ and $\beta_{2}=0.999$.
@@
Figure 5: Comparison of EM Accuracy between KV-Llama3.1 and Llama3.1 across different gold context positions. With varying gold context positions, KV-Llama3.1 illustrates consistent accuracies across datasets. However, Llama3.1 suffers from ‘lost in the middle’ problem, which can be resolved by KV-Fusion models.
@@ A.7 Position Agnostic reader evaluation on Contriever retrieved passages
Table 6: Accuracy comparison with other position-invariant methods on contriever-retrieved passages. KV-Fusion models achieve the highest accuracies across datasets except NQ top-5 case. Consistent with the results observed for DPR-retrieved passages in Table 1, KV-Fusion models show strong robustness to the inclusion of additional passages, while other methods experience a decline in performance as more passages are added. Notably, the strong performance on POPQA datasets highlights Contriever’s ability to excel on unseen datasets, which KV-Fusion models effectively leverage. This demonstrates that KV-Fusion models can achieve strong performance on different retrievers.

### v2

@@ 7 Conclusion
This paper presents KV-Fusion, a lightweight training scheme that addresses positional bias and improves the robustness of decoder-only models in the RAG pipeline. KV-Fusion trains language models to be context-order invariant by extracting and pre-filling KV caches with identical positional information and then training decoder-only models using these caches. The results highlight not only the robustness of KV-Fusion in handling a large number of input passages but also its position-invariant property. Our empirical evaluations on three open-domain datasets indicate that KV-Fusion can improve the performance and reliability of the RAG system.
@@ 8 Limitations
Second limitation is that our experiments are limited to single-hop question answering, where multi-step reasoning is not required. For example, datasets like HotpotQA Yang et al. (2018), and MuSiQue Trivedi et al. (2022) require multiple passages to derive answers. This work, however, focuses on single-hop question-answering datasets, making it difficult to assess the impact of KV-fusion in multi-hop datasets.
Third limitation is that this work does not fully explore the use of KV-cache for training LLMs. Recently, training LLMs by conditioning key-value caches have gained attention Sun et al. (2024) for its efficiency, though our approach remains underexplored in terms of language modeling. However, we present strong empirical results to solve the ‘Lost in the middle’ problem. We hope our work can facilitate future studies on utilizing key-value cache for training LLMs.
@@ A.1.1 Algorithm for KV-Fusion
11:     Update parameters of $\mathcal{D}_{\textrm{t}}$ with respect to $Loss$ via gradient descent
@@
To train $\mathcal{D}_{\textrm{t}}$ seamlessly with huggingface transformersWolf et al. (2020) and PyTorchPaszke et al. (2017), extracted KV-cache need to be reshaped to prefill the caches in $\mathcal{D}_{\textrm{t}}$. To this end, we implement res function down below, which can also process batch of instances.
def reshape_key_value_batches(cur_past_key_values, n_psgs):     """     Reshape key-value pairs in batches     """     new_key_cache = []     # Iterate through each key-value pair     for k, v in cur_past_key_values:         # Split keys and values into splits (split by instance)         k_splits, v_splits = torch.split(k, n_psgs, dim=0), torch.split(v, n_psgs, dim=0)         # Reshape and concatenate splits (reshape by instance)         k_re = torch.cat([torch.cat(torch.split(k_val, 1, dim=0), dim=2) for k_val in k_splits], dim=0)         v_re = torch.cat([torch.cat(torch.split(v_val, 1, dim=0), dim=2) for v_val in v_splits], dim=0)         # Append processed key-value pair         new_key_cache.append((k_re, v_re))     return tuple(new_key_cache)
@@ A.2.2 Dataset Statistics
NQ and TriviaQA are filtered versions provided by Karpukhin et al. 2020 under the CC BY-NC 4.0. We also use POPQA, available from the HuggingFace datasets [^3] https://huggingface.co/datasets/akariasai/PopQA under MIT License. These datasets are English open domain question answering datasets based on the December 2018 Wikipedia snapshot, preprocessed by Karpukhin et al. 2020.
@@ A.5 Hyperparameters for training
Table 5: Across all datasets, we utilize four A100 80 GB GPUs, with a batch size of 2 per device and a gradient accumulation of 8, consuming approximately 12 hours (48 GPU hours). The number of training passages is 20, consisting of one gold context and 19 negative contexts. Contexts are tokenized with a maximum length of 192 tokens using left padding; if a context exceeds this limit, tokens are truncated from the left. A random sample of 10,000 contexts from the NQ training set showed that 99.0% of contexts fit within this token limit.(Avg: 145 tokens, Std: 14 tokens) The maximum learning rate is set to $2\times 10^{-5}$, using a linear warmup and cosine decay. The warmup ratio is set to 5%. The AdamW optimizer(paged_adamw_32bit) is used with $\beta_{1}=0.9$ and $\beta_{2}=0.999$.
@@
Figure 5: Comparison of EM Accuracy between KV-Llama3.1 and Llama3.1 across different gold context positions. With varying gold context positions, KV-Llama3.1 illustrates consistent accuracies across datasets. However, Llama3.1 suffers from the ‘Lost in the middle’ problem, which can be resolved by KV-Fusion models.
@@ A.7 Position Agnostic reader evaluation on Contriever retrieved passages
Table 6: Accuracy comparison with other position-invariant methods on contriever-retrieved passages. KV-Fusion models achieve the highest accuracies across datasets except for NQ top-5 case. Consistent with the results observed for DPR-retrieved passages in Table 1, KV-Fusion models show strong robustness to the inclusion of additional passages while other methods experience a decline in performance as more passages are added. Notably, the strong performance on POPQA datasets highlights Contriever’s ability to excel on unseen datasets, which KV-Fusion models effectively leverage. This demonstrates that KV-Fusion models can achieve strong performance on different retrievers.